OpenStudy (anonymous):

[6.02] and [6.03] What is the value of the x variable in the solution to the following system of equations? 3x + y = 1 6x + 2y = −9 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (anonymous):

help please

OpenStudy (smokeysthename):

solve to find x.. use substitution. 3x+y=1 subtract y from both sides and subtract 1 from both sides. what will you get?

OpenStudy (anonymous):

im not sure i don't understand

OpenStudy (smokeysthename):

let me go slower. looking at the first equation only right now.. 3x+y=1 subtract y from both sides.

OpenStudy (anonymous):

okay

OpenStudy (smokeysthename):

you'll get 3x=-y+1 now subtract 1 from both sides. what will you get?